OpenStudy (anonymous):

Hi! I need help with some Hospitality homework with converting measurements into other portions. Anyone care to help me through it? I'm a bit' stumped as to what the first step should be. You are given the following recipe for egg noodle dough. Your supervisor asked you to use this recipe to make 20 pounds of egg noodle dough. Make the necessary adjustments to the following recipe to yield 20 pounds of egg noodle dough. Note this recipes yield is its total weight. Egg Noodles Bread Flour 5# Eggs, Whole 1# 8oz Salt 2oz Oil, Canola 1oz Water 1#

OpenStudy (anonymous):

I have steps that I wrote down for the method of answering these questions, I'm just not sure what the "new" or "old" yield is. My steps written are. 1. Convert the new and old yield to oz. 2. Divide the new yield by the old yield. 3. Get the number 4. Multiple new number by all ingredients. 5. Get all new numbers 6. Add all numbers together to check the answer Can anyone fill in the blanks for me? What am I missing out on here?

OpenStudy (anonymous):

Best thing to do first would get them all in ounces. There are 26 pounds in one ounce. Bread would be 5*16 Eggs (1*16) + 8 Salt 2 Oil 1 water 1*16 do the math and add all of them up and tell me what you get

OpenStudy (anonymous):

Sure! Give me just one moment. It just occurred to me that the # meant pounds. Duh, silly me. One sec!

OpenStudy (anonymous):

141 oz?

OpenStudy (anonymous):

5*16=80 1*16=16 16+8=24 =2 =1 1*16=16

OpenStudy (anonymous):

139? (◠﹏◠✿)

OpenStudy (anonymous):

80+24=104 104+2=106 106+1=107 107+16=123

OpenStudy (anonymous):

Which step is this? Or what is this leading from?

OpenStudy (anonymous):

this is converting all of the ingredients into ounces, so it is 123 ounces.

OpenStudy (anonymous):

you want to make 20 pounds so 20x16=320 thats 320 ounces, take that 320 and divide it by the original weight of ingredients and get 2.601 or 2.6. Now go back and multiply each ingredient by 2.6 and that is how much of each you need for 20 pounds of egg soup.

OpenStudy (anonymous):

I see now, the text was confusing me. I don't learn particularly well via text but my options aren't broad at the moment so I thank you for having the patience to assist me. Bread - 208 oz Whole Eggs - 62.4 Salt - 5.2 Oil - 2.6 Water - 41.6

OpenStudy (anonymous):

Or wait, it'd be each ingredient of the newer conversion?

OpenStudy (anonymous):

I have no idea. x-x I think that may be what I'm supposed to do right there?

OpenStudy (anonymous):

seems like you did it right. As long as they dont want it converted back to ounces and pounds, that should be you answer, in ounces.

OpenStudy (anonymous):

Thank you! You've been very helpful.

OpenStudy (anonymous):

Glad I could help.